OpenStudy (anonymous):

PLEASE IF ONLY SOMEONE CAN HELP ME WITH THIS!Shelly delivers the weekly local paper to neighborhoods in her town. House numbers are even on one side of the street and odd on the other. Shelly delivers an equal number of papers to both sides of the street. Although she always aims for the front doorstep, Shelly typically misses on three of the tosses on her route each week. Design and conduct a simulation to estimate the probability that next week, Shelly's three misses will all be at odd-numbered houses.

OpenStudy (anonymous):

a) Explain clearly your design of the simulation, including choice of probability tool and description of a single trial. b) Conduct the simulation with twenty trials and record the results. c) Calculate the experimental probability that all 3 of Shelly's missed papers will be at odd-numbered houses.

OpenStudy (anonymous):

do you think you can help me do this problem?

OpenStudy (anonymous):

Hello?

OpenStudy (dumbcow):

how many total houses are there?

OpenStudy (anonymous):

thats all the information i have from this but i think you can deside the number of houses

OpenStudy (anonymous):

the tool you could use is a standard die

OpenStudy (anonymous):

but then you have to conduct 20 trials

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Hint: You can set up the experiment using 3 coins to collect the data. Allow one side of the coin to represent Heads (evens) and one other side to represent Tails (odds).

OpenStudy (dumbcow):

sorry, well in that case, the probability of a miss being either odd/even is 1/2 thus probability all odd = 1/2*1/2*1/2 = 1/8

OpenStudy (dumbcow):

yeah a coin is good idea...or use single die and if you roll an even num that means the miss is an even house

OpenStudy (anonymous):

so it could be like odd, even , odd

OpenStudy (anonymous):

how do you do part b?

OpenStudy (dumbcow):

right, then record combination and repeat 20 times you should find that you only get all odd about 2 or 3 times out of 20

OpenStudy (anonymous):

so on each trial i would flip a coin three times

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

What did you get for the 20 trials?

OpenStudy (dumbcow):

haha im not doing the trials, you are what do you get :)

OpenStudy (anonymous):

thanks i finished it

OpenStudy (anonymous):

best response

OpenStudy (dumbcow):

:)